OpenStudy (anonymous):

Eliminate the "z" variable in the last two equations by finding opposite coefficients and adding. What is the resulting equation? x + 3y + 2z = -16 2x - y + 2z = -15 2x - 2y - z = 1

OpenStudy (anonymous):

Okay I got you give me a sec...

OpenStudy (anonymous):

Thank You!:)

OpenStudy (anonymous):

To eliminate a variable you multiply one of the equations by a number that will make the coefficient of the variable you want to get rid of have the opposite value as the coefficient of the same variable in the other equation. so, 2x-y+2z=-15 2x-2y-z=1 2x -y + 2z= -15 2(2x-2y-z)= 2(1) which is 2x -y + 2z= -15 4x-4y -2z = 2 You now add the like components of the equation (i.e. x with x, y with y, and constants with constants). This gives you: (2x+4x)+(-y-4y)+(2z-2z)=(-15+2) which simplifies to: 6x-5y+0=-13

OpenStudy (anonymous):

Thanks Cutie! :D

OpenStudy (anonymous):

Thanks ya (: